# Setting $\varepsilon$ is not the Issue in Differential Privacy

### v1

@@ 1 Introduction
The collapse of storage and data processing costs, along with digitization, has expanded the possibilities for machine learning (ML) and contributed to its recent successes. For instance, Large Language Models rely on trillions of tokens drawn from highly diverse sources villalobos2022will. Big data often involves sensitive information, making privacy a central challenge in trustworthy ML. Mitigating the risks to individuals whose data is included in training sets is crucial to comply with regulations, such as the GDPR gdpr or the California Consumer Privacy Act cpra, to avoid harming users, and to build long-term trust. However, efficiently protecting privacy while extracting value from the data remains a difficult task.
Differential Privacy (DP) originalDP has emerged as the gold standard in privacy-preserving machine learning. Its key idea is to provide a generic definition that imposes constraints on the output distribution of an algorithm – typically, the predictions of a machine learning model or the entire set of weights if the model is released publicly – such that, if these constraints are satisfied, no attacker can infer too much about the exact participants in the training dataset. Thus, DP quantifies, in the worst case, how much information about a single entity in the dataset can be leaked through its influence on the algorithm’s output, providing theoretical guarantees against all possible attacks. More precisely, an algorithm $\mathcal{A}$ is differentially private if, for all pairs of datasets $\mathcal{D}\sim\mathcal{D}^{\prime}$ and every measurable subset $\mathcal{S}\subset Z$, the following inequality holds:
@@
Many differentially private algorithms have been developed, spanning the full spectrum of machine learning, from statistical reporting and dimensionality reduction to synthetic data generation, reinforcement learning, gradient descent, and other optimization method dworksurvey; 9064731. Many approaches to ensuring DP rely on noise injection, which typically replaces the true value with a noised version, allowing randomness to propagate throughout the algorithm and into its outputs. However, this noise injection comes at a cost to the performance of the algorithm, typically resulting in less accurate predictions. For instance, the current state-of-the-art non-private baseline on ImageNet achieves around $90\%$, while a fully differentially private algorithm with a privacy budget of $\varepsilon=8$ achieves $39.2\%$ pmlr-v202-sander23b. Similarly, DP LLM tends to be five-year behind non-private version, but exists with $\varepsilon=2$ mckenna2025scalinglawsdifferentiallyprivate The trade-off between leaving an algorithm untouched and non-private, versus enforcing strong constraints that may destroy its utility, is governed by the privacy budget. Setting $\varepsilon$ is thus often the first step in running a differentially private algorithm. Once $\varepsilon$ is fixed, it is possible to roughly determine the noise scale at each step. The exact budget spent can then be computed more tightly during training using a numerical privacy accountant.
@@ 2 The many challenges of quantifying privacy
For instance, Latanya Sweeney demonstrated that, in the 1990 U.S. Census, 87.1% of respondents could be uniquely identified using just their ZIP code, birth date, and sex Sweeney2015OnlyYY. In particular, she famously mailed the health record of the governor of Massachusetts to him after he claimed that privacy was not jeopardized by the public release of anonymous health records histosweeney. Another well-known case is the Netflix Prize controversy, where researchers showed that by cross-referencing the Netflix dataset with another public dataset, the Internet Movie Database (IMDb), it was possible to link IMDb public profiles to complete movie histories in the Netflix data. This linkage could, in some cases, reveal sensitive information such as political or sexual orientation netflix, with varying probabilities of success depending on the number of rated movies in a user’s profile. To sum up, data can be either useful or perfectly anonymous, but never both brokenpromises, and our intuition of data leakage as a binary event does not reflect the reality of high-dimensional data, where the probability of re-identification increases with the number of features.
Privacy intuition is not only lacking when estimating the risk of re-identification, but also when deciding whether the release of information constitutes a privacy violation. Researchers have recently pointed out that even the release of public information – such as personal data disclosed as a contact in a patent – may be perceived as a privacy violation if memorized by a large language model (LLM) and surfaced in a completely different context, such as targeted advertising Tramr2022ConsiderationsFD. The fact that determining whether a flow of information is appropriate requires consideration of multiple parameters was formalized over twenty years ago by the Contextual Integrity framework nissenbaum2004privacy.
@@
In this section, we have so far recapped some difficulties in intuitively assessing privacy risks, and the fact that privacy expectations depend on several aspects of data flows. Another lesson from anonymization is that privacy risks are better captured by probabilities than by binary outcomes. However, human intuition about probabilities is also lacking. It has been shown that many people fail to recognize that $1/10$ is as large as $10/100$ ALONSO20031537. More specifically, in the context of probability estimation, even professionals such as doctors have been shown to violate basic principles of probability. The Linda Problem lindaproblem illustrates this: when asked to rank the likelihood of events involving a fictional feminist character named Linda, participants consistently judged the probability that she was an accountant to be lower than the probability that she was both an accountant and a feminist activist – even though the second event is a strict subset of the first, and thus necessarily less probable. Interestingly, Linda’s paradox may be partially explained by the fact that people do not estimate the absolute value of a probability, but rather how the probability is modified by the prior description of Linda. This conjunction fallacy is sometimes attributed to people estimating the plausibility of an event rather than its probability, effectively treating it as a conditional probability based on the given information. As this is very similar to the quantity estimated by differential privacy, it is not impossible that the privacy budget might actually align better with human intuition than other probability-based measures, even though the connection remains conjectural and, to our knowledge, has not been formally studied so far.
@@ 3 Selected advantages of Differential Privacy
Another clearly desirable property is adaptive composition, which is formally guaranteed by differential privacy. Adaptive composition ensures that if, after applying a first differentially private algorithm to a dataset, we apply a second algorithm to the same dataset – possibly using the output of the first – the final output remains differentially private, and we can compute the overall privacy budget pmlr-v37-kairouz15. This property is essential in settings where data can be reused, which is the default in most applications. In contrast, ad hoc privacy methods may fail under composition. For instance, applying $k$-anonymity sweeney2002Kanonymity twice to the same dataset can break $k$-anonymity and allow for unique re-identification. Similarly, methods like $l$-diversity and $t$-closeness, while intended to address some weaknesses of $k$-anonymity, are not closed under composition and may also degrade with repeated use or auxiliary information.
@@
In the case of randomized response ($RR$), the interpretation of the privacy budget is quite straightforward. For computing a histogram on $K$ possible outputs, $RR$ returns the true value with probability $e^{\varepsilon}/(K-1+e^{\varepsilon})$ and chooses uniformly at random among the K1K-1 other responses the rest of the time. In the case of $K=2$, the mitigation of privacy risk is so clear that randomized response was introduced before differential privacy as a method for surveying people on sensitive questions in 1965 randomizedresponse. Its implementation used two coin flips, ensuring $\varepsilon=\log(3)$. Although this privacy budget is greater than $1$ in that case (which is often taken as a bondary for small $\varepsilon$), it was still enough to establish trust in the mechanism. This illustrates that "all small $\varepsilon$ are alike" in the sense that users are likely to accept a modest risk even without a strict hard constraint on the privacy budget. A strength of differential privacy is that it provides worst-case guarantees, meaning that intuition based on low-dimensional cases is typically more pessimistic than the actual risk in high-dimensional settings. To illustrate this point, consider again the randomized response mechanism as $K$ grows, for instance, to the number of available emoji (3790). In this scenario, a randomized response with parameter $\log(3)$ returns an incorrect randomly selected response $99.92\%$ of the time. In comparison, achieving a $75\%$ chance of returning the true answer – similarly to the guarantee provided by $\varepsilon=\log(3)$ for binary randomized response – corresponds to $\varepsilon=\log(3\times 3789)\sim 9.33$, a value often seen as a very high privacy budget. This arises because the ratio of the probability of returning the true emoji compared to any other response increases drastically with $K$, so that while the ratio is large, the overall probability of returning any specific incorrect emoji remains small. In more complex scenarios, such as the output distribution of machine learning models, it is extremely hard to predict which measurable set $\mathcal{S}$ will maximize the ratio and whether it will be observed frequently enough to leak information in practice. Differential privacy addresses this issue by enforcing that, except for events small enough to be absorbed in $\delta$, the ratio is strictly bounded, which simplifies interpretation.
@@
The privacy budget can also be interpreted in a purely statistical setting, where one tests the null hypothesis that $x$ was part of the dataset against the alternative hypothesis that the dataset did not contain $x$. Differential privacy is then equivalent to requiring that any such hypothesis test must have either low significance – that is, a high rate of false positives (Type I errors) – or low power – that is, a high rate of false negatives (Type II errors). More precisely, when weighting one of these two errors by $e^{\varepsilon}$, their sum must always remain at least $1-\delta$ Wasserman01032010. More recent works have deepened the connection between hypothesis testing and differential privacy, allowing users with a statistics background to develop a strong intuition for privacy guarantees pmlr-v108-balle20a; hypstatforeps; Su2024ASV.
@@
Finally, an advantage of differential privacy is the growing literature that bridges the gap between the worst-case guarantees given by the privacy budget and the observed success of known attacks. This connection can be made through privacy auditing, where one empirically estimates, for fixed (possibly crafted) pairs of inputs, the actual probability of a set of outcomes $\mathcal{S}$, thus providing an empirical lower bound on the possible privacy budget. Auditing methods have recently improved in both precision and efficiency auditingtighter; Cebere2024TighterPA; steinke2023privacy; Mahloujifar2024AuditingFP. Numerous works have also connected various privacy budgets to the success rates of membership inference attacks and even full reconstruction attacks, covering a wide range of practical scenarios for translating privacy budgets into estimates of current risk quantifyeps; tablereidentification. Some studies have also explored in detail how to interpret the privacy budget in light of such empirical attacks inproceedings; nanayakkara2023epsilon; epswithattributeattacks. We give a visual explanation of the connection between hypothesis testing and empirical attacks in Figure 1(a)
@@ 4 The risk of focusing on minimizing $\varepsilon$ within DP research
For researchers already experienced in privacy-preserving machine learning, the motivations presented in the previous section – such as ensuring post-processing and composition properties – are likely already clear, and the paper may seem more relevant for newcomers. In this section, we argue that the controversy around setting the privacy budget also negatively impacts current research in differential privacy. In particular, we believe that the privacy budget exemplifies Goodhart’s Law goodhart1984problems, often summarized as: “When a measure becomes a target, it ceases to be a good measure.” We briefly recap the meaning of this law before describing how it affects the field, notably by encouraging weaker trust regimes and a tropism on small $\varepsilon$ results.
@@ 4.1 Decreasing $\varepsilon$ by unfair accounting
The fact that the signal of a measure tends to disappear when it is used as a metric predates Goodhart’s work, as it is an ubiquitous phenomenon found in diverse domains, from economics to medicine and education muller2019tyranny. This has been described from a statistician’s perspective by Desrosières in the context of French public policy Desrosires2005. One illustrative case is the treatment of unemployment statistics: when unemployment began to be seen as a politically relevant metric, it started to be gamed, and subsequent decreases often resulted more from new counting techniques excluding parts of the unemployed population than from real improvements in the job market Desrosires2014. More broadly, the sociology of quantification has shown that metrics used in decision-making processes often embed highly subjective choices and cannot be treated as complete or faithful representations of the underlying phenomena.
@@
In fact, the limitations of focusing on a single metric are widely documented in the machine learning community. For instance, the true significance of benchmark results is increasingly being questioned hardt2025emerging, and even seen as potentially harmful in the long term wang_against_2022. Metrics can be contradictory pmlr-v27-luxburg12a and may fail to predict real-world behavior when models are deployed NEURIPS2023_e304d374. A metric is typically a proxy for a much more complex objective, and this one-dimensional simplification is only useful as long as we do not attempt to optimize it at all costs. In such cases, improvements often result from artifacts in the evaluation process rather than from significant progress.
@@
Recent criticism of results that rely on pretraining with public datasets, allowing private data to be used only for fine-tuning is good example of this dangerous trend of artificially decreasing the privacy budgetTramr2022ConsiderationsFD. This strategy is less privacy-demanding and creates the impression that it is possible to train very large models under a small privacy budget with current methods. In particular, problematic pretraining on public data can be seen as a technique to game both accuracy and privacy metrics, which does not benefit privacy research. In many of the tested scenarios, there is overlap between the so-called private and public datasets (e.g., pretraining on CIFAR-100 or ImageNet and fine-tuning on CIFAR-10). This overlap undermines the meaningfulness of the reported accuracy, as these are not genuine transfer learning settings, high accuracy may not reflect effective private fine-tuning. It also illustrates how this approach encourages researchers to game the privacy budget: if private data is duplicated in public pretraining, it should be included in the privacy accounting to preserve end-to-end privacy guarantees. Finally, favoring this public pretraining setup because it enables small $\varepsilon$ values also tends to favor large models that cannot be run locally by users. This creates a centralization risk: users must trust a third party with access to their private data to produce outputs, potentially undermining the very privacy guarantees that differential privacy aims to provide.

### v2

@@ 1 Introduction
The collapse of storage and data processing costs, along with digitization, has expanded the possibilities for machine learning (ML) and contributed to its recent successes. For instance, Large Language Models rely on trillions of tokens drawn from highly diverse sources Villalobos et al. (2022). Big data often involves sensitive information, making privacy a central challenge in trustworthy ML. Mitigating the risks to individuals whose data is included in training sets is crucial to comply with regulations, such as the GDPR Parliament and of the European Union (2016) or the California Consumer Privacy Act Mactaggart (2020), to avoid harming users, and to build long-term trust. However, efficiently protecting privacy while extracting value from the data remains a difficult task.
Differential Privacy (DP) Dwork et al. (2006) has emerged as the gold standard in privacy-preserving machine learning. Its key idea is to provide a generic definition that imposes constraints on the output distribution of an algorithm – typically, the predictions of a machine learning model or the entire set of weights if the model is released publicly – such that, if these constraints are satisfied, no attacker can infer too much about the exact participants in the training dataset. Thus, DP quantifies, in the worst case, how much information about a single entity in the dataset can be leaked through its influence on the algorithm’s output, providing theoretical guarantees against all possible attacks. More precisely, an algorithm $\mathcal{A}$ is differentially private if, for all pairs of datasets $\mathcal{D}\sim\mathcal{D}^{\prime}$ and every measurable subset $\mathcal{S}\subset Z$, the following inequality holds:
@@
Many differentially private algorithms have been developed, spanning the full spectrum of machine learning, from statistical reporting and dimensionality reduction to synthetic data generation, reinforcement learning, gradient descent, and other optimization method Dwork (2008); Gong et al. (2020). Many approaches to ensuring DP rely on noise injection, which typically replaces the true value with a noised version, allowing randomness to propagate throughout the algorithm and into its outputs. However, this noise injection comes at a cost to the performance of the algorithm, typically resulting in less accurate predictions. For instance, the current state-of-the-art non-private baseline on ImageNet achieves around $90\%$, while a fully differentially private algorithm with a privacy budget of $\varepsilon=8$ achieves $39.2\%$ Sander et al. (2023). Similarly, DP LLM tends to be five-year behind non-private version, but exists with $\varepsilon=2$ McKenna et al. (2025) The trade-off between leaving an algorithm untouched and non-private, versus enforcing strong constraints that may destroy its utility, is governed by the privacy budget. Setting $\varepsilon$ is thus often the first step in running a differentially private algorithm. Once $\varepsilon$ is fixed, it is possible to roughly determine the noise scale at each step. The exact budget spent can then be computed more tightly during training using a numerical privacy accountant.
@@ 2 The many challenges of quantifying privacy
For instance, Latanya Sweeney demonstrated that, in the 1990 U.S. Census, 87.1% of respondents could be uniquely identified using just their ZIP code, birth date, and sex Sweeney (2015). In particular, she famously mailed the health record of the governor of Massachusetts to him after he claimed that privacy was not jeopardized by the public release of anonymous health records Sweeney (2000). Another well-known case is the Netflix Prize controversy, where researchers showed that by cross-referencing the Netflix dataset with another public dataset, the Internet Movie Database (IMDb), it was possible to link IMDb public profiles to complete movie histories in the Netflix data. This linkage could, in some cases, reveal sensitive information such as political or sexual orientation Narayanan and Shmatikov (2006), with varying probabilities of success depending on the number of rated movies in a user’s profile. To sum up, data can be either useful or perfectly anonymous, but never both Ohm (2009), and our intuition of data leakage as a binary event does not reflect the reality of high-dimensional data, where the probability of re-identification increases with the number of features.
Privacy intuition is not only lacking when estimating the risk of re-identification, but also when deciding whether the release of information constitutes a privacy violation. Researchers have recently pointed out that even the release of public information – such as personal data disclosed as a contact in a patent – may be perceived as a privacy violation if memorized by a large language model (LLM) and surfaced in a completely different context, such as targeted advertising Tramèr et al. (2022). The fact that determining whether a flow of information is appropriate requires consideration of multiple parameters was formalized over twenty years ago by the Contextual Integrity framework Nissenbaum (2004).
@@
In this section, we have so far recapped some difficulties in intuitively assessing privacy risks, and the fact that privacy expectations depend on several aspects of data flows. Another lesson from anonymization is that privacy risks are better captured by probabilities than by binary outcomes. However, human intuition about probabilities is also lacking. It has been shown that many people fail to recognize that $1/10$ is as large as $10/100$ Alonso and Fernández-Berrocal (2003). More specifically, in the context of probability estimation, even professionals such as doctors have been shown to violate basic principles of probability. The Linda Problem Tversky and Kahneman (1983) illustrates this: when asked to rank the likelihood of events involving a fictional feminist character named Linda, participants consistently judged the probability that she was an accountant to be lower than the probability that she was both an accountant and a feminist activist – even though the second event is a strict subset of the first, and thus necessarily less probable. Interestingly, Linda’s paradox may be partially explained by the fact that people do not estimate the absolute value of a probability, but rather how the probability is modified by the prior description of Linda. This conjunction fallacy is sometimes attributed to people estimating the plausibility of an event rather than its probability, effectively treating it as a conditional probability based on the given information. As this is very similar to the quantity estimated by differential privacy, it is not impossible that the privacy budget might actually align better with human intuition than other probability-based measures, even though the connection remains conjectural and, to our knowledge, has not been formally studied so far.
@@ 3 Selected advantages of Differential Privacy
Another clearly desirable property is adaptive composition, which is formally guaranteed by differential privacy. Adaptive composition ensures that if, after applying a first differentially private algorithm to a dataset, we apply a second algorithm to the same dataset – possibly using the output of the first – the final output remains differentially private, and we can compute the overall privacy budget Kairouz et al. (2015). This property is essential in settings where data can be reused, which is the default in most applications. In contrast, ad hoc privacy methods may fail under composition. For instance, applying $k$-anonymity Sweeney (2002) twice to the same dataset can break $k$-anonymity and allow for unique re-identification. Similarly, methods like $l$-diversity and $t$-closeness, while intended to address some weaknesses of $k$-anonymity, are not closed under composition and may also degrade with repeated use or auxiliary information.
@@
In the case of randomized response ($RR$), the interpretation of the privacy budget is quite straightforward. For computing a histogram on $K$ possible outputs, $RR$ returns the true value with probability $e^{\varepsilon}/(K-1+e^{\varepsilon})$ and chooses uniformly at random among the K1K-1 other responses the rest of the time. In the case of $K=2$, the mitigation of privacy risk is so clear that randomized response was introduced before differential privacy as a method for surveying people on sensitive questions in 1965 Warner (1965). Its implementation used two coin flips, ensuring $\varepsilon=\log(3)$. Although this privacy budget is greater than $1$ in that case (which is often taken as a bondary for small $\varepsilon$), it was still enough to establish trust in the mechanism. This illustrates that "all small $\varepsilon$ are alike" in the sense that users are likely to accept a modest risk even without a strict hard constraint on the privacy budget. A strength of differential privacy is that it provides worst-case guarantees, meaning that intuition based on low-dimensional cases is typically more pessimistic than the actual risk in high-dimensional settings. To illustrate this point, consider again the randomized response mechanism as $K$ grows, for instance, to the number of available emoji (3790). In this scenario, a randomized response with parameter $\log(3)$ returns an incorrect randomly selected response $99.92\%$ of the time. In comparison, achieving a $75\%$ chance of returning the true answer – similarly to the guarantee provided by $\varepsilon=\log(3)$ for binary randomized response – corresponds to $\varepsilon=\log(3\times 3789)\sim 9.33$, a value often seen as a very high privacy budget. This arises because the ratio of the probability of returning the true emoji compared to any other response increases drastically with $K$, so that while the ratio is large, the overall probability of returning any specific incorrect emoji remains small. In more complex scenarios, such as the output distribution of machine learning models, it is extremely hard to predict which measurable set $\mathcal{S}$ will maximize the ratio and whether it will be observed frequently enough to leak information in practice. Differential privacy addresses this issue by enforcing that, except for events small enough to be absorbed in $\delta$, the ratio is strictly bounded, which simplifies interpretation.
@@
The privacy budget can also be interpreted in a purely statistical setting, where one tests the null hypothesis that $x$ was part of the dataset against the alternative hypothesis that the dataset did not contain $x$. Differential privacy is then equivalent to requiring that any such hypothesis test must have either low significance – that is, a high rate of false positives (Type I errors) – or low power – that is, a high rate of false negatives (Type II errors). More precisely, when weighting one of these two errors by $e^{\varepsilon}$, their sum must always remain at least $1-\delta$ Wasserman and Zhou (2010). More recent works have deepened the connection between hypothesis testing and differential privacy, allowing users with a statistics background to develop a strong intuition for privacy guarantees Balle et al. (2020); Liu et al. (2019); Su (2024).
@@
Finally, an advantage of differential privacy is the growing literature that bridges the gap between the worst-case guarantees given by the privacy budget and the observed success of known attacks. This connection can be made through privacy auditing, where one empirically estimates, for fixed (possibly crafted) pairs of inputs, the actual probability of a set of outcomes $\mathcal{S}$, thus providing an empirical lower bound on the possible privacy budget. Auditing methods have recently improved in both precision and efficiency Nasr et al. (2023); Cebere et al. (2024); Steinke et al. (2023); Mahloujifar et al. (2025). Numerous works have also connected various privacy budgets to the success rates of membership inference attacks and even full reconstruction attacks, covering a wide range of practical scenarios for translating privacy budgets into estimates of current risk Lee and Clifton (2011); Jayaraman and Evans (2019). Some studies have also explored in detail how to interpret the privacy budget in light of such empirical attacks Balle et al. (2022); Nanayakkara et al. (2023); Pankova and Laud (2022). We give a visual explanation of the connection between hypothesis testing and empirical attacks in Figure 1(a)
@@ 4 The risk of focusing on minimizing $\varepsilon$ within DP research
For researchers already experienced in privacy-preserving machine learning, the motivations presented in the previous section – such as ensuring post-processing and composition properties – are likely already clear, and the paper may seem more relevant for newcomers. In this section, we argue that the controversy around setting the privacy budget also negatively impacts current research in differential privacy. In particular, we believe that the privacy budget exemplifies Goodhart’s Law Goodhart and Goodhart (1984), often summarized as: “When a measure becomes a target, it ceases to be a good measure.” We briefly recap the meaning of this law before describing how it affects the field, notably by encouraging weaker trust regimes and a tropism on small $\varepsilon$ results.
@@ 4.1 Decreasing $\varepsilon$ by unfair accounting
The fact that the signal of a measure tends to disappear when it is used as a metric predates Goodhart’s work, as it is an ubiquitous phenomenon found in diverse domains, from economics to medicine and education Muller (2019). This has been described from a statistician’s perspective by Desrosières in the context of French public policy Desrosières and Kott (2005). One illustrative case is the treatment of unemployment statistics: when unemployment began to be seen as a politically relevant metric, it started to be gamed, and subsequent decreases often resulted more from new counting techniques excluding parts of the unemployed population than from real improvements in the job market Desrosières (2014). More broadly, the sociology of quantification has shown that metrics used in decision-making processes often embed highly subjective choices and cannot be treated as complete or faithful representations of the underlying phenomena.
@@
In fact, the limitations of focusing on a single metric are widely documented in the machine learning community. For instance, the true significance of benchmark results is increasingly being questioned Hardt (2025), and even seen as potentially harmful in the long term Wang et al. (2022). Metrics can be contradictory von Luxburg et al. (2012) and may fail to predict real-world behavior when models are deployed Teney et al. (2023). A metric is typically a proxy for a much more complex objective, and this one-dimensional simplification is only useful as long as we do not attempt to optimize it at all costs. In such cases, improvements often result from artifacts in the evaluation process rather than from significant progress.
@@
Recent criticism of results that rely on pretraining with public datasets, allowing private data to be used only for fine-tuning is good example of this dangerous trend of artificially decreasing the privacy budgetTramèr et al. (2022). This strategy is less privacy-demanding and creates the impression that it is possible to train very large models under a small privacy budget with current methods. In particular, problematic pretraining on public data can be seen as a technique to game both accuracy and privacy metrics, which does not benefit privacy research. In many of the tested scenarios, there is overlap between the so-called private and public datasets (e.g., pretraining on CIFAR-100 or ImageNet and fine-tuning on CIFAR-10). This overlap undermines the meaningfulness of the reported accuracy, as these are not genuine transfer learning settings, high accuracy may not reflect effective private fine-tuning. It also illustrates how this approach encourages researchers to game the privacy budget: if private data is duplicated in public pretraining, it should be included in the privacy accounting to preserve end-to-end privacy guarantees. Finally, favoring this public pretraining setup because it enables small $\varepsilon$ values also tends to favor large models that cannot be run locally by users. This creates a centralization risk: users must trust a third party with access to their private data to produce outputs, potentially undermining the very privacy guarantees that differential privacy aims to provide.